# Dynamic compensation for pump-induced frequency shift in Kerr-cat qubit initialization

### v1

@@ Abstract
Figure 1: (a) Bloch sphere representation of the Kerr-cat qubit encoding space. Here, |±$\left|\pm\alpha\right\rangle$ denotes the coherent states and $\left|C_{\alpha}^{k}\right\rangle=(\left|\alpha\right\rangle+k\left|-\alpha\right\rangle)/\sqrt{2}$ for $k=+,-,+\mathrm{i},-\mathrm{i}$. (b) Energy levels of the system as a function of $\Delta(0)$, the mode detuning at $t=0$ in Eq. 1 with no external pump. (c) Solid lines: evolution of the energy levels for an ideal case with $\Delta\equiv 0$ when ramping up the pump amplitude $\epsilon_{2}$ to a final value of $\epsilon_{\mathrm{2,KC}}$. Dashed lines: the removal of the degeneracy due to the pump-induced frequency shift (PIFS). (d) Evolution of the energy levels for a static compensation with a fixed pump detuning $\Delta_{\mathrm{as}}(t)\equiv 1.5K$. At $\epsilon_{2}=0$, $\Delta(0)=1.5K$, corresponding to the right dashed line in (b). In (c) and (d), the purple arrows indicate adiabatic evolution paths that map the vacuum state to $\left|C_{\alpha}^{+}\right\rangle$ and to $\left|\psi_{\mathrm{e,1}}^{+}\right\rangle\approx[\hat{D}(\alpha)+\hat{D}(-\alpha)]\left|1\right\rangle$, respectively.
@@
$\hat{H}_{\mathrm{cat}}=\Delta(t)a^{\dagger}a-K\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+\varepsilon_{2}(t)\hat{a}^{\dagger 2}+\varepsilon_{2}^{*}(t)\hat{a}^{2},$ (1)
where $\hat{a}^{\dagger}(\hat{a})$ is the creation (annihilation) operator of the mode, $\Delta$ is the relative frequency detuning of the mode, $K$ is the Kerr nonlinearity, and $\varepsilon_{2}$ is the pump amplitude. In an ideal experimental configuration with $\Delta=0$, the bosonic mode has a stabilized two-dimensional Hilbert space spanned by |±$\left|\pm\alpha\right\rangle$, i.e., the coherent states with an average photon number $|\alpha|^{2}=\varepsilon_{2}/K$. Consequently, the superposition states $\left|C_{\alpha}^{k}\right\rangle=(\left|\alpha\right\rangle+k\left|-\alpha\right\rangle)/\sqrt{2}$ for $k=+,-,+\mathrm{i},-\mathrm{i}$ are also stabilized states in the mode. Figure 1(a) presents the Bloch sphere of the Kerr-cat qubit by treating $\left|C_{\alpha}^{\pm}\right\rangle$ as the computational basis. The Kerr-cat qubit has the advantage of being bias-preserving, with the states along the X-axis having an enhanced lifetime that increases exponentially with the Kerr-cat size $|\alpha|^{2}$ Puri et al. (2017, 2020).
@@
The initialization or encoding of the Kerr-cat qubit can be realized by ramping up the amplitude of the squeezing pump $\varepsilon_{2}$. As shown by the solid lines in Fig. 1(c), the eigenstates of the Hamiltonian evolve from the Fock state basis $\left|0\right\rangle$, $\left|1\right\rangle$} to the computational basis of the Kerr-cat qubit following the traces as $\varepsilon_{2}$ increases. However, in practical experiments there are parasitic effects on the mode when applying pumps to it. Specifically, we encounter a dynamic PIFS $\Delta_{\mathrm{PI}}(t)=\gamma\varepsilon_{2}^{2}(t)$ with a coefficient $\gamma$. Consequently, the corresponding detuning varies with the pump as $\Delta=\Delta_{\mathrm{as}}-\Delta_{\mathrm{PI}}$, where $\Delta_{\mathrm{as}}=\omega_{\mathrm{a}}-\omega_{\mathrm{s}}/2$ is the detuning between the external squeezing pump ($\omega_{\mathrm{s}}$) and the bare mode ($\omega_{\mathrm{a}}$). We define the pump amplitude at the end of the ramping-up time $t_{\mathrm{up}}$ as $|\varepsilon_{2}(t_{\mathrm{up}})|=\varepsilon_{\mathrm{2,KC}}$ and the detuning $\Delta_{\mathrm{as}}(t_{\mathrm{up}})=\Delta_{\mathrm{as,KC}}$. Then, $\Delta_{\mathrm{PI}}(0)=0$ and $\Delta_{\mathrm{PI}}(t_{\mathrm{up}})=\Delta_{\mathrm{PI,KC}}=\gamma\varepsilon^{2}_{\mathrm{2,KC}}$. Here the subscript KC denotes the parameters for preserving the Kerr-cat qubit.

### v2

@@ Abstract
Figure 1: (a) Bloch sphere representation of the Kerr-cat qubit encoding space. Here, |±$\left|\pm\alpha\right\rangle$ denotes the coherent states and $\left|C_{\alpha}^{k}\right\rangle=(\left|\alpha\right\rangle+k\left|-\alpha\right\rangle)/\sqrt{2}$ for $k=+,-,+\mathrm{i},-\mathrm{i}$. (b) Energy levels of the system as a function of $\Delta(0)$, the detuning at $t=0$ in Eq. 1 with no external pump. (c) Solid lines: evolution of the energy levels for an ideal case with $\Delta\equiv 0$ when ramping up the pump amplitude $\epsilon_{2}$ to a final value of $\epsilon_{\mathrm{2,KC}}$. Dashed lines: the removal of the degeneracy due to the pump-induced frequency shift (PIFS). (d) Evolution of the energy levels for a static compensation with a fixed pump detuning $\Delta_{\mathrm{as}}(t)\equiv 1.5K$. At $\epsilon_{2}=0$, $\Delta(0)=1.5K$, corresponding to the right dashed line in (b). In (c) and (d), the purple arrows indicate adiabatic evolution paths that map the vacuum state to $\left|C_{\alpha}^{+}\right\rangle$ and to $\left|\psi_{\mathrm{e,1}}^{+}\right\rangle\approx[\hat{D}(\alpha)+\hat{D}(-\alpha)]\left|1\right\rangle$, respectively.
@@
$\hat{H}_{\mathrm{cat}}=\Delta(t)a^{\dagger}a-K\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+\varepsilon_{2}(t)(\hat{a}^{\dagger 2}+\hat{a}^{2}),$ (1)
where $\hat{a}^{\dagger}(\hat{a})$ is the creation (annihilation) operator of the mode, $\Delta$ is the relative frequency detuning, $K$ is the Kerr nonlinearity, and $\varepsilon_{2}$ is real and represents the pump amplitude. In an ideal experimental configuration with $\Delta=0$, the bosonic mode has a stabilized two-dimensional Hilbert space spanned by |±$\left|\pm\alpha\right\rangle$, i.e., the coherent states with an average photon number $|\alpha|^{2}=\varepsilon_{2}/K$. Consequently, the superposition states $\left|C_{\alpha}^{k}\right\rangle=(\left|\alpha\right\rangle+k\left|-\alpha\right\rangle)/\sqrt{2}$ for $k=+,-,+\mathrm{i},-\mathrm{i}$ are also stabilized states in the mode. Figure 1(a) presents the Bloch sphere of the Kerr-cat qubit by treating $\left|C_{\alpha}^{\pm}\right\rangle$ as the computational basis. The Kerr-cat qubit has the advantage of being bias-preserving, with the states along the X-axis having an enhanced lifetime that increases exponentially with the Kerr-cat size $|\alpha|^{2}$ Puri et al. (2017, 2020).
@@
The initialization or encoding of the Kerr-cat qubit can be realized by ramping up the amplitude of the squeezing pump $\varepsilon_{2}$. As shown by the solid lines in Fig. 1(c), the eigenstates of the Hamiltonian evolve from the Fock state basis $\left|0\right\rangle$, $\left|1\right\rangle$} to the computational basis of the Kerr-cat qubit following the traces as $\varepsilon_{2}$ increases. However, in practical experiments there are parasitic effects on the mode when applying pumps to it. Specifically, we encounter a dynamic PIFS $\Delta_{\mathrm{PI}}(t)=\gamma\varepsilon_{2}^{2}(t)$ with a coefficient $\gamma$. Consequently, the corresponding detuning varies with the pump as $\Delta=\Delta_{\mathrm{as}}-\Delta_{\mathrm{PI}}$, where $\Delta_{\mathrm{as}}=\omega_{\mathrm{a}}-\omega_{\mathrm{s}}/2$ is the detuning between the external squeezing pump ($\omega_{\mathrm{s}}$) and the bare mode ($\omega_{\mathrm{a}}$). We define the pump amplitude at the end of the ramping-up time $t_{\mathrm{up}}$ as $\varepsilon_{2}(t_{\mathrm{up}})=\varepsilon_{\mathrm{2,KC}}$ and the detuning $\Delta_{\mathrm{as}}(t_{\mathrm{up}})=\Delta_{\mathrm{as,KC}}$. Then, $\Delta_{\mathrm{PI}}(0)=0$ and $\Delta_{\mathrm{PI}}(t_{\mathrm{up}})=\Delta_{\mathrm{PI,KC}}=\gamma\varepsilon^{2}_{\mathrm{2,KC}}$. Here the subscript KC denotes the parameters for preserving the Kerr-cat qubit.